### v3

# Position: Membership Inference Attack Should Move On to Distributional Statistics for Distilled Generative Models

Abstract
To detect unauthorized data usage in training large-scale generative models (e.g., ChatGPT or Midjourney), membership inference attacks (MIA) have proven effective in distinguishing a single training instance (i.e., a member) from a single non-training instance (i.e., a non-member). This success is mainly credited to a memorization effect: models tend to perform better (e.g., yield higher log-likelihood) on a member than a non-member. However, we find that standard MIAs fail against distilled generative models (i.e., student models) that are increasingly deployed in practice for efficiency (e.g., ChatGPT 4o-mini). Trained exclusively on data generated from a large-scale generative model (i.e., a teacher model), the student model lacks direct exposure to any members (i.e., the teacher model’s original training data), thus nullifying the memorization effect that standard MIAs rely on. This finding reveals a serious privacy loophole, where generation-service providers could deploy a student model whose teacher was potentially trained on unauthorized data, yet claim the deployed model is “clean” because it was not directly trained on such data. Hence, are distilled models inherently unauditable for upstream privacy violations, and should we discard them when we care about privacy? We contend no, as we uncover a memory chain connecting the student and teacher’s member data: the distribution of student-generated data aligns more closely with the distribution of the teacher’s members than with non-members, which makes it possible to detect unauthorized data usage even when direct instance-level memorization is absent. This leads us to posit that MIAs on distilled generative models should shift from instance-level scores to distribution-level statistics. We further propose three principles of distribution-based MIAs for detecting unauthorized training data through distilled generative models, and validate our position through an exemplar framework. We lastly discuss the implications of our position.
1 Introduction
Recent advances in large-scale generative models have set new standards for synthesizing high-quality content across modalities, such as images (Ho et al., 2020) and languages (Brown et al., 2020). These advances have been quickly translated into successful commercialization through online services such as ChatGPT (Roumeliotis and Tselikas, 2023) and Midjourney (Zierock and Jungblut, 2023). However, the extensive datasets required to train these models often contain sensitive information from individuals who may not have explicitly consented to the use of their data for model development (Liu et al., 2024b). This concern is particularly pressing given the widespread adoption of large language models (LLMs) (Floridi and Chiriatti, 2020) and diffusion models (Ho et al., 2020; Song et al., 2023) in commercial applications. In this context, instance-level membership inference attacks (I-MIAs) (Carlini et al., 2022), designed to detect whether a single instance is used in training, offer a valuable auditing mechanism for detecting potential privacy violations and unauthorized data usage.
Figure 1: Model distillation raises privacy concerns. MIA can detect unauthorized data in the teacher model but fails when only the student model is available. The reason behind this failure is that the student model is trained with the teacher’s outputs only, rather than the original data (i.e., members). More importantly, based on the failure of MIAs here, generation-service providers can only publish the student models as a service, to bypass unauthorized data detection and claim they do not use any unauthorized data for training.
The success of I-MIAs mainly relies on a phenomenon in training such large-scale generative models: well-trained models tend to overfit to their training set111 Note that, for I-MIAs used in classifiers, overfitting is not the only way to help distinguish between members and non-members. For example, Carlini et al. (2022) successfully identified member data by analyzing model behavior signals, such as the output confidence distribution and gradient norms, even when a classifier does not overfit the data. Moreover, it does not mean that the overfitting is a golden rule that always makes I-MIAs successful. When the large-scale models are trained with extremely large datasets (e.g., ChatGPT), I-MIAs could be meaningless (Zhang et al., 2024). We will discuss this in detail in Section 3.1., exhibiting different behaviors between training instances (i.e., called members) and test instances (i.e., called non-members) (Yeom et al., 2018), which is also known as the memorization effect (Yeom et al., 2018). Based on this phenomenon, many I-MIAs consider using training objectives as criteria to distinguish between members and non-members. For example, for diffusion models, Duan et al. (2023) use estimation errors during the denoising process as the member-detection criterion; for large language models, Ye et al. (2024) and Shejwalkar and Houmansadr (2021) use log-likelihood for the same purpose. Recent studies (Carlini et al., 2023; Hu and Pang, 2023; Duan et al., 2023) have investigated these criteria extensively, hypothesizing that generative models can memorize their training instances.
However, recent generative model-based I-MIAs do not consider another important step in deploying powerful large-scale generative models: distillation, and, in this paper, we find that existing I-MIAs fail to find members when facing a distilled generative model. The requirement of distilled generative models is based on an urgent need of efficient deployment of powerful generative models, as deploying generative models on a scale requires substantial computational resources, often necessitating hundreds or even thousands of expensive GPUs (Hu et al., 2024). Specifically, during distillation, we will learn lightweight generative models (a.k.a. student models) with the data generated by a large-scale generative model (a.k.a. teacher models). Well-known examples include ChatGPT 4o-mini (Prazeres et al., 2025) and DeepSeek-V3 (Liu et al., 2024a). As such, model distillation enables a two-tier deployment strategy for generation service providers: original teacher models focus on training student models, while lightweight distilled models serve end-users.
Based on the nature of model distillation: student models only learn from teacher models’ outputs, it can be seen that the two-tier deployment strategy introduces a critical security caveat: since the student model does not learn from the training data of teachers (i.e., members), as Fig. 1 shows, the memorization effect, a key to the success of I-MIAs, is no longer true for student models and members. Namely, existing I-MIAs probably do not work when we can only access the distilled models. To further verify the above argument, we provide more empirical evidence for this security caveat in Sec. 3.1. We find that, while I-MIAs effectively identify training data in teacher models (Fig. 3(a)), they consistently fail with student models (Fig. 3(b)), implying that student models retain insufficient membership information at the instance level.
This finding poses a serious privacy issue: generation-service providers can only publish the student models as a service, to bypass unauthorized data detection and claim they do not use any unauthorized data for training, causing distilled generative models to naturally have privacy concerns when they are deployed. For example, to reduce costs, a model provider may crawl high-quality social media data without licensing to train an LLM (Part A in Fig. 1). To avoid copyright liability, they distill a student model from the teacher’s outputs and release it publicly (Part B). Since I-MIAs can only access the student model, which contains only instance-level signals of the teacher’s outputs, they fail to trace the unauthorized data use, leaving social media content unprotected.
The failure of I-MIA raises a question: Should we discard the distilled generative models when we care about privacy? The answer is negative, as we discover a distribution-based memory chain between a student-teacher pair: student-generated data exhibit a significantly stronger distributional alignment with teacher’s members than non-members, making it possible to determine if a student model has knowledge from unauthorized data. Specifically, in Sec. 4.1, we use maximum mean discrepancy (MMD) to compare student-generated samples with both teacher’s member and non-member data distributions (Fig. 3(c)). Through repeated random subset samplings, we observe a consistent statistical pattern–distances to non-member data concentrate at higher values than to member data, suggesting that the student preserves statistical signals exhibiting stronger alignment with the teacher’s member distribution than non-member distributions, despite the failure of I-MIAs. Thus, to reliably audit the privacy violation of distilled generative models, we posit the following statement in the field of generative models:
Position: Membership Inference Attacks (MIAs) for distilled generative models should shift from instance-level scores to distribution-level statistics.
Based on the position above, we suggest that auditing upstream privacy violation risks on distilled generative models should be evaluated based on the distribution of data instances rather than individual instances due to the discovered memory chain between a student-teacher pair, namely distribution-level MIA (D-MIA). In Sec. 5, we further establish three principles that D-MIAs should follow to maximize their effectiveness. Following these principles, we build an exemplar framework to evaluate whether a distribution of outputs from the distilled model more closely resembles the statistical properties of the members of the teacher model compared to non-members, and verify the effectiveness of the framework with preliminary experiments. We finally discuss the broader implications of privacy regulation and responsible AI deployment that D-MIA leads to, highlighting both opportunities and challenges in the evolving landscape of generative model auditing (Sec. 6).
2 Preliminary: MIAs for Generative Models
We first introduce some basic concepts and relevant literature in this section. For readers who are familiar with MIAs for generative models, it is safe to skip this section. MIAs evaluate whether a specific instance is used during model training. Let ${\mathcal{X}}$ be a data space and ${\mathcal{D}}^{\rm mem}\subset{\mathcal{X}}$ be a member set used to train a generative model $G:{\mathcal{Z}}\to{\mathcal{X}}$ that transforms noises sampled from a latent distribution ${\mathbf{z}}\sim p({\mathbf{z}})$ into synthetic data ${\mathbf{x}}=f_{g}({\mathbf{z}})\in{\mathcal{X}}$. Given a query sample ${\bm{x}}_{q}\in{\mathcal{X}}$, an MIA constructs a binary classifier ${\mathcal{A}}:{\mathcal{X}}\times G\to\{0,1\}$ that predicts the membership attribution of ${\mathbf{x}}_{q}$ as ${\mathcal{A}}({\bm{x}}_{q},G)=1$ if ${\bm{x}}_{q}\in{\mathcal{D}}^{\rm mem}$, and 00 otherwise. Existing MIAs use instance-level scores to distinguish members from non-members, a strategy we term instance-level MIA (I-MIA), including reference-based and intrinsic-based methods that are introduced below.
2.1 Brief introduction for instance-level MIAs
Reference-based I-MIAs use carefully constructed reference models to compute a score that can distinguish members from non-members. Given a target generative model $G$, one needs to construct $n$ architecture-similar or identical reference models $\{G_{i}^{\rm ref}\}_{i=1}^{n}$, leading to two complementary sets of models for a query sample ${\bm{x}}_{q}$,
${\mathcal{M}}_{1}=\{G_{i}^{\rm ref}:{\bm{x}}_{q}\in{\mathcal{D}}^{\rm mem}_{i}\}\text{ and }{\mathcal{M}}_{0}=\{G_{i}^{\rm ref}:{\bm{x}}_{q}\notin{\mathcal{D}}^{\rm mem}_{i}\},$
where ${\mathcal{D}}^{\rm mem}_{i}$ denotes the training dataset of the $i$-th reference model. The membership inference decision is then based on the difference between the target model $\phi(G,{\bm{x}})$ and these groups. For example, ${\mathcal{A}}({\bm{x}},G)=1$ if a difference metric $\Delta({\bm{x}},G)>\tau$ and 00 otherwise, where $\Delta({\bm{x}},G)\triangleq\mathrm{sim}(\phi(G,{\bm{x}}),{\mathcal{M}}_{1})-\mathrm{sim}(\phi(G,{\bm{x}}),{\mathcal{M}}_{0})$, $\mathrm{sim}(\cdot,\cdot)$ is a function to measure the similarity between two models, and $\tau$ is the decision threshold. However, this approach faces practical infeasibility for computationally intensive generative models like diffusion models, as training $n\geq 1$ reference models would increase the attack cost.
Intrinsic-based I-MIAs directly leverage the statistical gaps that emerge from target model training. At their core, these attacks exploit a fundamental memorization tendency of generative models $G:{\mathcal{Z}}\to{\mathcal{X}}$, i.e., the target model behaves differently between member instances ${\mathcal{D}}^{\rm mem}$ and non-member instances ${\mathcal{D}}^{\rm non}$, quantified as $\Delta({\mathbf{x}},G)=\mathbb{E}_{{\bm{x}}\sim{\mathcal{D}}^{\rm mem}}\left[{\mathcal{L}}({\bm{x}};G)\right]-\mathbb{E}_{{\bm{x}}^{\prime}\sim{\mathcal{D}}^{\rm non}}\left[{\mathcal{L}}({\bm{x}}^{\prime};G)\right]<0$, where ${\mathcal{L}}({\bm{x}};G)$ is normally the minimal distance between ${\bm{x}}$ and the data generated by $G$. This statistical gap may manifest differently across generative architectures, leading to model-specific attack strategies. GAN-Leak (Chen et al., 2020), for example, targets MIA on generative adversarial networks (Goodfellow et al., 2020) by reconstructing target images through latent optimization, solving ${\mathcal{L}}_{\rm GANLeak}({\mathbf{x}})=\min_{{\bm{z}}\in{\mathcal{Z}}}\|{\mathbf{x}}-G({\bm{z}})\|_{2}^{2}$ where members ${\bm{x}}\sim{\mathcal{D}}^{\rm mem}$ often show lower reconstruction errors.
2.2 How I-MIAs are applied to large-scale generative models?
Large-scale generative model in text domain-LLMs (Guo et al., 2025; Touvron et al., 2023) are auto-regressive transformers (Tang et al., 2024) trained on massive text corpora to model the log-likelihood of each next token given its preceding context. Several MIA strategies have been developed for LLMs. PCA (Ye et al., 2024) detects membership by comparing the target input with a synthetic version generated by word-swapping. A large log-likelihood gap indicates membership. The Min-K(Shokri et al., 2017) computes the average log-likelihood of the $k$ tokens with the least confidence in a given input. Typically, member data exhibit higher log-likelihoods for these tokens compared to non-members. Recall (Xie et al., 2024) introduces a non-member prefix to condition the model and quantifies the resulting change in log-likelihood. If the log-likelihood changes significantly, the data is inferred to be a member sample.
Large-scale generative model in vision domain-Diffusion models (Ho et al., 2020; Song et al., 2020), which operate through a forward process $q({\bm{x}}_{t}|{\bm{x}}_{0})$ that progressively adds Gaussian noise $\epsilon\sim{\mathcal{N}}(0,{\mathbf{I}})$ to the input ${\bm{x}}_{0}$ and a reverse noise removal process $p_{G}({\bm{x}}_{0}|{\bm{x}}_{t})$ that reconstructs the original ${\bm{x}}_{0}$ over time steps $t\in[0,T]$. Several attacks have been proposed in this context and have been evaluated on a class of diffusion models, the denoising diffusion probabilistic model (DDPM) (Ho et al., 2020). According to DDPM objective loss $\mathcal{L}_{\text{MSE}}=\left\|\epsilon_{G}(x_{t},t)-\epsilon\right\|^{2}$, SecMI (Duan et al., 2023) performs an attack by monitoring this loss across different time-steps $t$. ReDiffuse (Li et al., 2024) investigates reconstruction stability under noise perturbations, based on the observation that member data tend to yield more consistent reconstructions. GSA (Pang et al., 2023) examines gradient dynamics during model retraining and finds that member data typically induces smaller parameter updates.
3 Instance-level MIAs Are Not Suitable for Distilled Generative Models
In this section, we demonstrate that I-MIAs are not capable of detecting members of a large-scale generative model when we can only access its distilled student model.
3.1 Large-scale generative models challenge the effectiveness of I-MIAs
I-MIAs face challenges when attacking LLMs. Although I-MIA remains widely used for large-scale generative models, it is unreliable, specifically for LLMs (Dong et al., 2024; Ye et al., 2024). This is because extensive training on massive corpora and substantial model capacity of LLMs would lead to similar behaviors between individual members and non-members exploited by I-MIAs (Ye et al., 2024). When processing an input, LLMs consistently produce high-confidence outputs regardless of whether it is part of training data, reducing the discriminative power of instance-level metrics (Dong et al., 2024). Moreover, Zhang et al. (2024) argues that I-MIAs on LLMs suffer from unboundable false positive rates (FPR)—a critical metric for evaluating the validity of MIAs when used as evidence to allege the use of unauthorized data (Carlini et al., 2022; Zhang et al., 2024). The reason is that LLMs are trained with massive web-scale corpora, such that there are no ‘true’ non-members (as LLMs indirectly learn too much data (Zhang et al., 2024)).
I-MIAs face challenges when attacking diffusion models. In the case of generative models in the visual domain—diffusion models, existing MIAs have successfully identified members in non-distilled settings (Duan et al., 2023; Li et al., 2024). However, there are still issues for both reference-based MIAs and intrinsic-based MIAs. First, reference-based MIAs are impractical for diffusion models due to (1) the high cost of training a reference model (Ho et al., 2020), and (2) the impossibility of recovering the original input noise, which is sampled randomly during training (Ho et al., 2020; Kim et al., 2023).
Figure 2: Model distillation for large-scale generative models. Diffusion models and LLMs rely on synthetic data produced by teacher models rather than the original training data to train student models. Thus, student models do not see the original training sets.
Second, several intrinsic-based MIAs have shown promising results on DDPMs (Ho et al., 2020). Since the DDPM objective (described in Sec. 2.2) encourages accurate reconstruction of the forward exact noise vector $\epsilon$, the model tends to overfit the denoising trajectories, resulting in lower reconstruction loss for members compared to non-members. Because DDPMs are too time-consuming for deployment (as they require thousands of denoising steps (Ho et al., 2020)), to reduce the number of denoising steps, more effective consistency-based models have been proposed to learn a globally consistent denoising trajectory. These efficient methods are achieved by incorporating a KL smoothing term into the objective function (Kim et al., 2023; Song et al., 2023; Karras et al., 2022). However, since the KL smoothing term guides the model toward a consistent denoising trajectory across samples rather than memorizing sample-specific trajectories, it reduces the loss discrepancy between member and non-member data and may weaken the effectiveness of intrinsic-based MIAs.
3.2 Distillation exacerbates I-MIA failures in generative models
Figure 3: Comparison of I-MIAs on teacher model EDM (Karras et al., 2022) and student model DMD (Yin et al., 2024) using: (a) ReDiffuse (Li et al., 2024) reveals membership signals in EDM via distinct reconstruction and re-noising losses between members and non-members on AFHQv2. (b) Applied to DMD, ReDiffuse fails to separate member from non-member instances. (c) Student outputs show stronger distributional alignment with member data when evaluated as instance sets via MMD (Gretton et al., 2012).
Distilled models for online generation service. Although large-scale generative models can produce high-quality text and images, the billion-scale parameters of LLMs and iterative denoise steps of diffusion models lead to high inference latency (Touvron et al., 2023; Song et al., 2020), presenting challenges for online deployment of generation services. Knowledge distillation (Hinton, 2015) offers a compelling solution, enabling the creation of smaller “student” models that learn to mimic the output of larger “teacher” models and achieve gains in inference efficiency without compromising the generation quality (Hsieh et al., 2023; Gu et al., 2023). Examples abound. In LLMs, methods like MiniLLM (Gu et al., 2023) train student models on prompts paired with the teacher’s generated responses. In diffusion models, distillation efforts (Yin et al., 2024; Luo et al., 2024; Song et al., 2023) aim to reduce the multi-step denoising procedure into a single generation step, drastically speeding up inference with a focus on maintaining image quality (Meng et al., 2023; Duan et al., 2023; Luo et al., 2024). This suggests an evolving paradigm of how large-scale generation services are deployed: end-users are increasingly likely to interact with efficient student models, while the computationally intensive teacher models will be dedicated to training these efficient alternatives offline.
However, state-of-the-art distillation practices implement strict separation: student models learn exclusively from teacher-generated data, with no direct access to the teacher’s training dataset (see Fig. 2). Consequently, the distillation procedure introduces a barrier between the student model and the teacher’s original training instances (members). We therefore argue that this separation prevents the student from forming instance-level memories of the teacher’s members. As a result,
The memorization effect, which I-MIAs rely on, will not apply to distilled generative models.
Empirical evidence. We thus investigate the impact of distillation on I-MIAs. Fig. 3(b) shows that the student model’s reconstruction pattern exhibits no statistically significant differences between member and non-member images. We confirm this using four I-MIA methods on a teacher diffusion model (EDM (Karras et al., 2022)) and its student models (DMD (Yin et al., 2024) and Diff-Instruct (Luo et al., 2024)). Detailed setup is in App. A. Moreover, Tab. 1 shows that I-MIAs achieve a success rate higher than random guessing when applied to the teacher model, but perform no better than random guessing on student models. Since student models do not directly fit the teacher’s member data, they may not preserve the instance-level behavioral signature that I-MIAs typically exploit. Thus, model distillation, primarily developed for efficiency though, provides an inherent defense against major I-MIAs (Shejwalkar and Houmansadr, 2021; Tang et al., 2022), causing I-MIAs to fail against distilled models.
4 Does Distillation Really Eliminate Membership Information?
Although the distilled generative models are trained without seeing any members of the teacher model, their learning process could be influenced by these members: the student trains on data generated by the teacher, who aims to approximate the data distribution from which its members were drawn. We therefore conjecture the following:
The student, in learning to mimic the teacher’s output distribution, might indirectly learn a distribution closer to the teacher’s member distribution than to a distribution of non-members.
We call this potential propagation of distributional characteristics the memory chain between student and member data, and empirically confirm its existence.
4.1 Distributional signals survive distillation
Table 1: Average performance of MIAs on three generative models: EDM (teacher model), DMD, and Diff-Instruct (distilled models). Metrics include ASR, AUC, and TPR@FPR=0.05. See App. C for detail.
Model Dataset ASR AUC TPR@FPR=0.05
EDM CIFAR10 0.596 0.610 0.070
FFHQ 0.584 0.590 0.083
AFHQv2 0.704 0.724 0.230
DMD CIFAR10 0.515 0.509 0.050
FFHQ 0.515 0.503 0.047
AFHQv2 0.526 0.520 0.063
Diff-Instruct CIFAR10 0.508 0.505 0.043
FFHQ 0.508 0.509 0.050
AFHQv2 0.509 0.508 0.053
We first define a distributional membership signal in the context of distillation as the tendency for a collection of samples generated by the student model to inherit the teacher’s distributional bias and, in some statistical sense, appear closer to the distribution induced by the member data.
The key question now becomes does a student trained on teacher-generated data indeed preserve such a distributional membership signal? We test it empirically. Consider three datasets: student-generated ${\mathcal{D}}^{\rm gen}$, teacher’s member data ${\mathcal{D}}^{\rm mem}$ and disjoint non-member data ${\mathcal{D}}^{\rm non}$, we evaluate the distance between ${\mathcal{D}}^{\rm gen}$ and ${\mathcal{D}}^{\rm mem}$ against the distance between ${\mathcal{D}}^{\rm gen}$ and ${\mathcal{D}}^{\rm non}$ across multiple experimental trials for statistical robustness, randomly sampling subsets $\tilde{{\mathcal{D}}}^{\rm gen}$, $\tilde{{\mathcal{D}}}^{\rm mem}$, and $\tilde{{\mathcal{D}}}^{\rm non}$ from their respective datasets each trial. We adopt maximum mean discrepancy (MMD) (Gretton et al., 2012), a kernel-based distance measure between probability distributions, to quantify distributional similarities between paired subsets, namely (i) $\tilde{{\mathcal{D}}}^{\rm gen}$ and $\tilde{{\mathcal{D}}}^{\rm mem}$, and (ii) $\tilde{{\mathcal{D}}}^{\rm gen}$ and $\tilde{{\mathcal{D}}}^{\rm non}$. We observe a pattern across repeated trials: The MMD values of (i) cluster at lower magnitudes compared to those of (ii), indicating that the student-generated data aligns more closely with the teacher’s member distribution from its non-member counterpart (Fig. 3(c)). This way we confirm that distribution-level statistics (e.g., distribution discrepancy) can identify the teacher’s membership information undetected at the instance level, even through a distilled student model. Namely, distributional membership signals survive under the distillation procedure, and the memory chain between a student-teacher pair exists.
4.2 Distributional statistics enables reliable membership inference with LLMs
The merits of focusing on distributional statistics for membership inference are not limited to distilled models. This perspective provides a more robust approach in general, as models scale and training data become more complex, particularly in contexts of MIAs for LLMs, where I-MIAs fall short.
Distributional statistics amplify instance-level membership signals. Ye et al. (2024) show that LLMs exhibit characteristic uncertainty patterns across local neighborhoods of training data, revealing membership information invisible to single-sample analyses. Similarly, Dong et al. (2024) confirm that membership information lies in set-level probability distributions rather than individual confidence scores. Such findings align with our observations: membership information manifests collectively more obviously than individual instances, requiring aggregate statistical measures to detect.
Distributional information alleviates the FPR issue. The shift to distributional analysis also addresses a critical flaw plaguing conventional I-MIAs on LLMs (Zhang et al., 2024). As discussed in Sec. 3.1, I-MIAs applied to LLMs often yield an unacceptably high FPR, since computing FPR reliably requires access to “true” non-member data which, model has never truly seen, even indirectly. Given that LLMs are trained on massive web-scale corpora (Liu et al., 2024b), which may contain paraphrases, summaries, or citations of virtually any public data, guaranteeing such pristine non-exposure is practically impossible. Distributional approaches, however, are more resilient to this issue by focusing on aggregate statistical properties of data, such as token frequency distributions or specific sampling behaviors. Even if models see scattered references to non-member datasets, such exposure proves insufficient to reproduce the complete statistical signal of a dataset (Choi et al., 2025). Unless substantial portions of a dataset appear in training, models cannot exhibit the coherent distributional patterns that distributional MIAs detect, enabling reliable membership inference even in contaminated evaluation.
5 Membership Inference Attacks Should Move On to Distributional Statistics
Our findings in Sec. 4 confirm the existence of a memory link that links data generated by a student model back to its teacher’s original member data, This, coupled with the greater reliability of distribution-level statistics over instance-level scores, leads us to a central argument: privacy audits for distilled generative models should shift to distributional statistics. This section outlines three guiding principles for such distribution-based MIAs and introduces D-MIA, a pilot example framework embodying these principles, to verify their practical value. Still, we acknowledge the limitations of D-MIA in its current iteration and call for future research efforts.
5.1 Guiding principles for distributional membership inference
Reflecting on the challenges of I-MIAs (Sec. 3.1) and the persistence of distributional signals through distillation (Sec. 4), we propose that effective MIAs, especially where instance-level cues are weak (as in model distillation), should be built upon the following principles: (1) Set-based analysis: Distributional measures are inherently more stable and reveal more pronounced signals when computed over sets of data points rather than isolated instances (Sec . 4.2). MIAs should leverage this by analyzing collections of samples to enhance their statistical power. (2) Distributional comparison: The insight from Sec. 4.1 is that a student’s output distribution is often statistically closer to its teacher’s member distribution than to non-member distributions. Effective MIAs should therefore quantify this statistical divergence to uncover traces of memorization. (3) Discriminative signals focus: Effective MIAs should target signals whose discriminative power between members and non-members becomes more salient at the distributional level, while retaining their resilience even through transformations, such as model distillation.
5.2 D-MIA: an exemplar framework for auditing distilled generative models
To illustrate how these principles can be put into practice, we showcase a pilot implementation called D-MIA, which considers the following problem setup. Let $G_{\rm T}$ be a teacher generative model pre-trained on a private member dataset, denoted by $\mathcal{D}^{{\rm mem}}$. Except for them, samples are collectively denoted by $\mathcal{D}^{{\rm non}}$ otherwise. We assume access to a student generative model $G_{\rm S}$, trained only on the generated data from $G_{\rm T}$. D-MIA enables set-based analysis by task design. Given a candidate dataset $\mathcal{D}^{\text{can}}$, the goal is to determine whether $\mathcal{D}^{\text{can}}$ overlaps with the teacher’s member dataset $\mathcal{D}^{\text{mem}}$.
For distributional comparison, D-MIA quantifies and examines the relative relationship between two quantities: 1) the distributional distance between candidate dataset ${\mathcal{D}}^{\rm can}$ and student-generated dataset ${\mathcal{D}}^{\rm gen}$, and 2) the distributional distance between known non-member data ${\mathcal{D}}^{\rm non}$ and student-generated data ${\mathcal{D}}^{\rm gen}$.
To focus on discriminative signals, D-MIA uses a two-stage approach. During training, it optimizes a deep kernel MMD-based measure (Liu et al., 2020) to distinguish member from non-member data. This involves training a kernel that maximizes the distributional separation between known members ${\mathcal{D}}^{\rm mem}$ and non-members ${\mathcal{D}}^{\rm non}$ relative to the student-generated data ${\mathcal{D}}^{\rm gen}$. In evaluation, this learned metric determines whether a particular candidate set ${\mathcal{D}}^{\rm can}$ is statistically more similar to ${\mathcal{D}}^{\rm gen}$ than to ${\mathcal{D}}^{\rm non}$. If so, ${\mathcal{D}}^{\rm can}$ is inferred to likely contain member data (Further technical details are in App. B.2).
Table 2: ASR and AUC results of D-MIA against baselines I-MIA methods SecMI, and ReDiffuse on distilled models across CIFAR10, FFHQ, and AFHQv2. Rows are color-coded to represent member data proportions: 100%, 50%, and 30%. See Tab. 8 in App. C for the TPR@FPR=0.05 results.
Dataset (Member %) DMD Diff-Instruct
D-MIA SecMI ReDiffuse D-MIA SecMI ReDiffuse
ASR AUC ASR AUC ASR AUC ASR AUC ASR AUC ASR AUC
CIFAR10 (100%) 0.98 0.99 0.60 0.55 0.66 0.66 1.0 1.0 0.65 0.54 0.62 0.62
CIFAR10 (50%) 0.98 0.99 0.59 0.52 0.60 0.60 1.0 1.0 0.59 0.53 0.60 0.60
CIFAR10 (30%) 0.92 0.97 0.53 0.43 0.60 0.59 1.0 1.0 0.53 0.47 0.52 0.55
FFHQ (100%) 1.0 1.0 0.60 0.56 0.56 0.56 1.0 1.0 0.57 0.56 0.78 0.81
FFHQ (50%) 0.99 0.99 0.56 0.54 0.54 0.49 1.0 1.0 0.55 0.52 0.65 0.63
FFHQ (30%) 0.98 0.99 0.56 0.49 0.54 0.48 1.0 1.0 0.55 0.51 0.62 0.59
AFHQv2 (100%) 1.0 1.0 0.61 0.60 0.69 0.71 1.0 1.0 0.56 0.53 0.64 0.62
AFHQv2 (50%) 1.0 1.0 0.59 0.54 0.64 0.61 1.0 1.0 0.53 0.48 0.57 0.52
AFHQv2 (30%) 1.0 1.0 0.56 0.56 0.60 0.61 1.0 1.0 0.48 0.50 0.55 0.50
5.3 Empirical support
D-MIA is effective against distilled generative models. Tab. 2 shows that D-MIA can successfully detect membership across various distilled models and datasets, even when candidate sets are mixtures of member and non-member data. For example, against DMD (a state-of-the-art distillation technique), D-MIA achieves near-perfect success rates (ASR $\approx 100\%$) across three datasets, significantly outperforming baselines. D-MIA maintains high ASR ($\approx 92\%$) on CIFAR10 with mixed candidate datasets, while baselines falter to near-random guessing if only $30\%$ of the candidate sets are members. This confirms D-MIA as a reliable framework for practical scenarios where candidate sets often have unknown compositions of member and non-member data.
D-MIA can quantify dataset composition. Beyond a binary decision, D-MIA’s output scores correlate positively with the portion of member data in a candidate set (See Appendix Fig. 5). Scores tend towards 1 for entirely member-comprised sets and decrease towards 0.5 as member presence diminishes, suggesting a new role for D-MIA in more granular privacy leakage analysis.
5.4 Alternative view: contexts where I-MIAs remain relevant
While the shift to distributional statistics, exemplified by D-MIA, addresses key challenges in auditing distilled generative models, there are practical considerations that stress why I-MIAs–despite their limitations–remain useful for certain privacy auditing scenarios.
When candidate data is scarce. Distributional methods like D-MIA require a reasonable number of samples in a candidate set to estimate distributional statistics reliably (e.g., using MMD). In practice, data subjects (e.g., artists) may have only a limited collection of personal records–perhaps fewer than 10 pieces or even a single artwork–when they seek an audit to determine if their data was used to train a generative model. As Tab. 3 shows, D-MIA’s discriminative power degrades when candidate set sizes lower down (see App. B.5 for setup details). On the contrary, I-MIAs are not limited in this case as they probe model behavior at the sample level.
Resource considerations. All MIAs need reference data (known members and non-members). However, distributional comparisons might implicitly encourage the use of larger reference datasets to ensure robust statistical estimation. This could increase data storage and management overhead, potentially conflicting with data minimization principles under regulations like GDPR (Mondschein and Monda, 2019). Though future work may mitigate D-MIA’s resource demands via compression or more efficient samplings, such solutions remain speculative; I-MIAs already function under milder assumptions.
Granularity of privacy protection. Often, the core privacy concern revolves around a specific data point. An artist, for instance, typically wants to know if one particular artwork was used–the legal question concerning that single item’s membership, not just a general pattern across their entire body of work. I-MIAs, when successful, can provide this precise attribution.
6 Implications: Lessons from Distillation for Broader MIA
Table 3: ASR and AUC results of D-MIA evaluated on DMD under varying non-member and candidate dataset sizes. In each configuration, we equally split ${\mathcal{D}}^{\rm non}$ for kernel training and MIA evaluation. All metrics decrease as ${\mathcal{D}}^{\rm non}$ and ${\mathcal{D}}^{\rm can}$ lower down.
$|{\mathcal{D}}^{\rm non}|$ $|{\mathcal{D}}^{\rm can}|$ ASR AUC TPR@FPR=0.05
(5000+10000) 5000 0.98 0.99 0.96
(2000+4000) 2000 0.94 0.97 0.88
(600+1200) 600 0.83 0.78 0.70
(300+600) 300 0.75 0.58 0.52
Redefining membership and privacy harm in model lifecycles. D-MIAs compel us to expand the notion of “membership”. Instead of solely referring to “a specific instance the model was trained on”, membership can also signify “the statistical property that the model learns from the dataset”.
This expanded view is important for understanding privacy harm when models learn from transformed data–like student models trained on teacher-generated data–that still carry the statistical bias of the original, potentially sensitive member data. This concern, for example, may not be limited to whether a model can reproduce a specific photo, but whether it has learned to mimic a creator’s distinct artistic style or absorbed societal biases from a text corpus, even without memorizing exact data points. This shift highlights the need for a more comprehensive discussion about privacy risks and what constitutes unauthorized data use in the complex lifecycles of modern AI models.
Strengthening audits for model provenance and countering “model laundering”. As AI models and their training data become valuable assets, concerns about “model laundering”–the practice of obscuring the use of unauthorized data through techniques like distillation–are likely to grow. If a teacher model ($G_{\rm T}$) was trained on copyright-protected or sensitive data, and a company deploys a student model ($G_{\rm S}$) claiming that it was trained only on legitimate, teacher-generated data. I-MIAs on $G_{\rm S}$ would likely find no evidence of the original data misuse in training $G_{\rm T}$. In contrast, a successful distributional MIA against $G_{\rm S}$ could reveal that its learned data distribution closely matches that of the potentially problematic dataset used for $G_{\rm T}$, thus providing a crucial tool for auditing provenance and detecting such attempts to conceal unauthorized data use.
Distributional MIAs are more secure auditing tools. Recall that I-MIAs seek to identify individual data instances, which raises a security dilemma as well: tools designed to audit for privacy leakage by spotting specific training examples could, in the wrong hands, be abused to extract those same sensitive data. D-MIAs mitigate this tension by shifting focus from individual samples to candidate sets, evaluating alignment with the training distribution. D-MIAs assess whether a candidate dataset, as a whole, aligns with the training distribution’s characteristics. They can confirm significant data overlap without pinpointing which specific samples were members. Consequently, even if an attacker understands the distributional MIA mechanism, they cannot directly use the attack to determine the membership status of individual data points. From a privacy perspective, this dataset-based evaluation offers a new auditing paradigm with built-in privacy safeguards.
7 Final Remarks
This paper revealed that conventional I-MIAs falter against distilled models, creating a privacy loophole where original training data provenance can be obscured. We argued that the path forward requires a shift: MIAs should move beyond instance-level scores to embrace distribution-level statistics. We showed that while distillation may erase the direct memorization effect of individual members, the statistical pattern of member data is retained and can be exploited to perform effective MIA. We have accordingly proposed guiding principles for implementing distributional MIAs and verified their efficacy. We believe the transition to distributional MIAs is essential for effective privacy auditing and call for continued research to develop robust tools based distributional statistics, examine the privacy risks inherent in model distillation and beyond, and reassess the definitions of membership and privacy harm in generative AI.
Acknowledgments
ZSY, and FL are supported by the Australian Research Council (ARC) with grant number DE240101089, and FL is also supported by ARC with grant number LP240100101, DP230101540 and the NSF&CSIRO Responsible AI program with grant number 2303037. This research was also supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.
Appendix
Appendix A Setup for Diffusion Model and Distilled Models
The training configurations for EDM, DMD, and DI are shown in Tab. 4. The specific model architectures will be released in the upcoming official code. For each dataset, half of the data is randomly selected for training EDM, while the remaining half is used as non-member data. EDM generates 100,000 samples to distill DMD and DI models. During the distillation process, the models do not access the training data of EDM.
Table 4: Training configurations for different models (EDM, DMD, and DI) across datasets (CIFAR10, FFHQ, and AFHQv2), including GPU setups, batch sizes, training times, and learning rates.
Model Dataset GPU Batch size Training Time Learning Rate
EDM CIFAR10 1 $\times$ NVIDIA A100 128 5-00:00:00 0.001
FFHQ 4 $\times$ NVIDIA A100 256 5-00:00:00 0.0002
AFHQv2 2 $\times$ NVIDIA A100 128 5-00:00:00 0.0002
DMD CIFAR10 1 $\times$ NVIDIA A100 128 4-00:00:00 0.00005
FFHQ 1 $\times$ NVIDIA A100 64 4-00:00:00 0.00005
AFHQv2 1 $\times$ NVIDIA A100 64 4-00:00:00 0.00005
DI CIFAR10 1 $\times$ NVIDIA A100 128 3-00:00:00 0.00001
FFHQ 1 $\times$ NVIDIA A100 64 3-00:00:00 0.0001
AFHQv2 1 $\times$ NVIDIA A100 64 2-00:00:00 0.0001
Appendix B Details of D-MIA Framework
B.1 Preliminaries: MMD and Deep-kernel MMD
This section briefly summarizes the basic knowledge of MMD (Gretton et al., 2012) and its extension, deep-kernel MMD (Liu et al., 2020). MMD finds common application in areas such as domain adaptation and the evaluation of generative models, where assessing distributional alignment is important. In this paper, these concepts are used in quantifying the distributional differences in Section 4 and Section 5. We refer interested readers to the original papers for complete details therein.
Maximum Mean Discrepancy (MMD), proposed by  Gretton et al. (2012), is a statistical tool for measuring the distance between two Borel probability measures, say ${\mathbb{P}}$ and ${\mathbb{Q}}$, defined on a separable metric space ${\mathcal{X}}\subseteq{\mathbb{R}}^{d}$. Consider independent random variables $X,X^{\prime}\sim{\mathbb{P}}$ and $Y,Y^{\prime}\sim{\mathbb{Q}}$. The squared MMD between ${\mathbb{P}}$ and ${\mathbb{Q}}$ in a Reproducing Kernel Hilbert Space ${\mathbb{H}}_{k}$, induced by a kernel function $k:{\mathcal{X}}\times{\mathcal{X}}\to{\mathbb{R}}$, is defined as:
$\operatorname{MMD}^{2}(\mathbb{P},\mathbb{Q};k)=\mathbb{E}[k(X,X^{\prime})]+\mathbb{E}[k(Y,Y^{\prime})]-2\mathbb{E}[k(X,Y)].$
If $k$ is a characteristic kernel (e.g., Gaussian), then $\operatorname{MMD}^{2}(\mathbb{P},\mathbb{Q};k)=0$ if and only if ${\mathbb{P}}={\mathbb{Q}}$.
Empirical MMD. In practice, the true distributions ${\mathbb{P}}$ and ${\mathbb{Q}}$ are often unknown, and we rely on finite samples drawn from them. Given i.i.d. samples ${\mathcal{S}}_{X}=\{x_{i}\}_{i=1}^{n}$ from ${\mathbb{P}}$ and ${\mathcal{S}}_{Y}=\{y_{j}\}_{i=1}^{m}$ from ${\mathbb{Q}}$, an unbiased U-statistic estimator for $\operatorname{MMD}^{2}$ is
$\widehat{\operatorname{MMD}}_{u}^{2}({\mathcal{S}}_{X},{\mathcal{S}}_{Y};k)=\frac{1}{n(n-1)}\sum_{i\neq l}^{n}k(x_{i},x_{l})+\frac{1}{m(m-1)}\sum_{j\neq p}^{m}k(y_{i},y_{j})-\frac{2}{nm}\sum_{i=1}^{n}\sum_{j=1}^{n}k(x_{i},y_{j}).$ (1)
Deep-kernel MMD. Traditional MMD uses a fixed, pre-defined kernel, which may lead to limited expressiveness when the kernel is not suitable for the task at hand. To address this, Liu et al. (2020) propose to learn a task-relevant representation $\theta_{\omega}:{\mathcal{X}}\to{\mathcal{Z}}$ using a neural network parameterized by $w$. The MMD can then be computed in this learned feature space ${\mathcal{Z}}$. As such, the goal of deep-kernel MMD is to find a representation $\theta_{\omega}$ that maximizes the MMD, thereby increasing the test power to detect differences between ${\mathbb{P}}$ and ${\mathbb{Q}}$.
Following (Liu et al., 2020), let ${\mathcal{S}}_{X}=\{x_{i}\}_{i=1}^{n}$ and ${\mathcal{S}}_{Y}=\{y_{j}\}_{i=1}^{n}$ be samples from ${\mathbb{P}}$ and ${\mathbb{Q}}$ (assuming equal sample sizes $n$ for simplicity). The empirical estimate for deep-kernel MMD, using a U-statistic, is
$\widehat{\operatorname{MMD}}^{2}_{u}({\mathcal{S}}_{X},{\mathcal{S}}_{Y};k_{\omega}):=\frac{1}{n(n-1)}\sum_{i\neq j}H_{ij},$ (2)
where $H_{ij}$ is the kernel of the U-statistic, defined as
$H_{ij}:=k_{\omega}(x_{i},x_{j})+k_{\omega}(y_{i},y_{j})-k_{\omega}(x_{i},y_{j})-k_{\omega}(y_{i},x_{j}).$ (3)
Note that the kernel $k_{\omega}(\cdot,\cdot)$ itself is a composite function incorporating the learned features:
$k_{\omega}(a,b)=\left[\left(1-\epsilon\right)\,k_{\rm base}\left(\theta_{\omega}(a),\theta_{\omega}(b)\right)+\epsilon\right]\cdot q_{\rm base}\left(a,b\right).$ (4)
Here, $w$ is the feature extractor network (e.g., a multi-layer perceptron). $k_{\rm base}(\cdot,\cdot)$ is a base characteristic kernel (e.g., Gaussian) applied to the learned features $\theta_{\omega}(a)$ and $\theta_{\omega}(b)$. In addition, $q_{\rm base}(\cdot,\cdot)$ is typically another characteristic kernel on the original inputs, acting as a sample-pair weighting function that adjusts the influence of each pair in the kernel computation based on their importance or relevance. The small constant $\epsilon\in[0,1]$ helps to ensure $k_{w}$ remains characteristic (Liu et al., 2020).
Optimizing a deep-kernel MMD. When optimizing the parameters $w$ of the feature network $\theta_{\omega}$ to maximize the MMD estimate, it is often normalized by an estimate of its standard deviation to improve numerical stability and test power. The objective function is thus:
$\max_{\omega}{\mathcal{L}}({\omega})=\frac{\widehat{\operatorname{MMD}}^{2}_{u}({\mathcal{S}}_{X},{\mathcal{S}}_{Y};k_{\omega})}{\sigma\left(\widehat{\operatorname{MMD}}^{2}_{u}({\mathcal{S}}_{X},{\mathcal{S}}_{Y};k_{\omega})\right)},$ (5)
where $\sigma(\cdot)$ denotes the standard deviation of the MMD estimator. Since the true variance $\sigma^{2}$ is generally unknown, we estimate it using a regularized estimator $\hat{\sigma}_{\lambda}^{2}$, given by:
$\hat{\sigma}^{2}_{\lambda}=\frac{4}{n^{3}}\sum_{i=1}^{n}\left(\sum_{j=1}^{n}H_{ij}\right)^{2}-\frac{4}{n^{4}}\left(\sum_{i=1}^{n}\sum_{j=1}^{n}H_{ij}\right)^{2}+\lambda,$ (6)
where $\lambda$ is a constant to avoid division by zero. The optimization of $w$ is typically performed using stochastic gradient ascent on ${\mathcal{L}}(w)$.
B.2 D-MIA illustration
We now introduce the details of D-MIA, a pilot implementation that adheres to the guiding principles (Sec. 5) and verify the effectiveness of distributional statistics in performing MIA.
Problem setting. Since D-MIA operates with sets of data instances, it is noteworthy to mention that we consider a new problem setup than conventional I-MIAs.
Let $G_{\rm T}:{\mathcal{Z}}\to{\mathcal{X}}$ be a teacher generative model pre-trained on a private member dataset ${\mathcal{D}}^{\rm mem}=\{{\bm{x}}_{i}\}_{i=1}^{N}$, where ${\bm{x}}_{i}\sim{\mathbb{P}}^{\rm mem}$, ${\mathbb{P}}^{\rm mem}$ is the member data distribution. We have access to a distilled student generative model $G_{\rm S}$ that mimics $G_{\rm T}$’s behavior, trained using synthetic samples $\{G_{\rm T}({\bm{z}}_{j})\}_{j=1}^{M}$ with noises ${\bm{z}}_{j}\sim{\mathbb{P}}_{{\mathcal{Z}}}$, the distribution over the latent space (e.g., standard Gaussian). In D-MIAs, we consider set-based prediction: given a candidate dataset ${\mathcal{D}}^{\rm can}=\{{\bm{x}}_{j}^{\prime}\}_{j=1}^{N}$, the task is to infer if ${\mathcal{D}}^{\rm can}\cap{\mathcal{D}}^{\rm mem}=\varnothing$, i.e., contains member instances.
D-MIA requires two reference datasets: (1) a non-member set ${\mathcal{D}}^{\rm non}=\{{\bm{x}}_{k}^{\prime\prime}\}_{k=1}^{N}$ of public instances ${\bm{x}}_{k}^{\prime\prime}\nsim{\mathbb{P}}^{\rm mem}$ and (2) an anchor set ${\mathcal{D}}^{\rm anc}=\{{\bm{x}}_{l}^{*}\}_{l=1}^{L}$ (e.g., generated by $G_{\rm S}$) used to facilitate distributional comparison. Moreover, since private member data is typically inaccessible, we propose to construct a proxy member set $\widetilde{{\mathcal{D}}}^{\rm mem}=\{G_{\rm S}({\bm{z}}_{j})\}_{j=1}^{N}$ to approximate ${\mathbb{P}}^{\rm mem}$. At its core, D-MIA aims to detect whether ${\mathcal{D}}^{\rm can}$ aligns more closely with $\widetilde{{\mathcal{D}}}^{\rm mem}$ or ${\mathcal{D}}^{\rm non}$ through relative distributional discrepancy thresholding.
Training a deep-kernel MMD. We first optimize a data-adaptive kernel $k_{\omega}$, parameterized by deep neural nets $\omega$ (Liu et al., 2020) to maximize the separation between $\tilde{{\mathcal{D}}}^{\rm mem}$ and ${\mathcal{D}}^{\rm non}$ in the feature space. For $\tilde{{\mathcal{D}}}^{\rm mem}$, ${\mathcal{D}}^{\rm non}$ and ${\mathcal{D}}^{\rm anc}$, we perform mini-batch training and randomly sample subsets from each dataset, e.g., ${\mathcal{B}}^{\rm anc}=\{{\bm{x}}_{b}^{*}\overset{{\rm i.i.d}}{\sim}{\mathcal{D}}^{\rm anc}\}_{b=1}^{B}$, with respect to the optimization objective ${\mathcal{L}}(\omega)$ defined as
${\mathcal{L}}(\omega)=\underbrace{\left[\widehat{\rm MMD}^{2}_{u}({\mathcal{B}}^{\rm anc},\tilde{{\mathcal{B}}}^{\rm mem};k_{\omega})\right]}_{\text{member discrepancy}}-\underbrace{\left[\widehat{\rm MMD}^{2}_{u}({\mathcal{B}}^{\rm anc},{\mathcal{B}}^{\rm non};k_{\omega})\right]}_{\text{non-member discrepancy}}.$
Doing so amplifies the MMD values between non-members and the anchor distribution while minimizing them for member-like distributions. See Alg. 1 for details.
Detecting membership. In this step, we aim to determine whether ${\mathcal{D}}^{\rm can}\cap{\mathcal{D}}^{\rm mem}=\varnothing$, by computing two MMD statistics using the trained kernel $k_{\omega}$: $M_{1}^{(t)}\triangleq\widehat{\rm MMD}^{2}_{u}({\mathcal{B}}^{\rm anc},{\mathcal{B}}^{\rm can};k_{\omega})$ and $M_{2}^{(t)}\triangleq\widehat{\rm MMD}^{2}_{u}({\mathcal{B}}^{\rm anc},{\mathcal{B}}^{\rm non};k_{\omega})$ over $T$ Bernoulli trials. The membership is indicated per trial via ${\mathbb{I}}^{(t)}=\mathbbm{1}(M_{1}<M_{2})$, and the aggregate membership probability is estimated by $p^{\rm mem}=\frac{1}{T}\sum_{t}{\mathbb{I}}^{(t)}$ (details are in Alg. 2).
Ensembling multiple kernels. To mitigate the variance from finite-sample MMD estimates (Chérief-Abdellatif and Alquier, 2022), we aggregate predictions across $m$ independently trained kernels $\{k_{\omega}^{(i)}\}_{i=1}^{m}$. For each kernel, we compute $p^{\rm mem})_{(i)}$ over $n$ Bernoulli trials as with Alg. 2. We apply a final decision threshold $\tau$ to the ensemble mean $\bar{p}^{\rm mem}=\frac{1}{m}\sum_{i}p^{\rm mem}_{(i)}$, declaring membership of ${\mathcal{D}}^{\rm can}$ if $\bar{p}^{\rm mem}>\tau$. See Alg. 3 for detailed illustrations.
Figure 4: Overview of our two-phase MMD-based D-MIA framework, consisting of (1) deep-kernel MMD training phase (top left) and (2) detecting the Candidate Dataset phase (bottom left). We also propose a kernel ensemble strategy to improve detection robustness (right).
B.3 Experimental setup
Dataset and victim Models. We empirically evaluate D-MIA on state-of-the-art distilled generative models, DMD (Yin et al., 2024) and Diff-Instruct (Luo et al., 2024) on commonly studied MIA benchmarks, CIFAR10 (Krizhevsky et al., 2010), FFHQ (Karras, 2019), and AFHQv2 (Choi et al., 2020). See detailed setup of victim models in App. A
Baseline settings. D-MIA differs from existing MIA methods and attack targets. To ensure fairness, we adapt existing methods to the D-MIA setting for experimentation. Specifically, we apply existing MIA methods to each data point in the dataset to compute a loss-based result. Then we compute the mean loss result of all data points in the dataset. We randomly sample 50 candidate datasets (with replacement) and 50 non-member datasets (with replacement) and calculate the mean loss for each dataset. Then, we empirically determine an optimal threshold to distinguish between the loss means of candidate datasets and non-member datasets. Under this setting, we use SecMI and ReDiffuse as baseline methods for comparison.
Evaluation settings. Before the experiment, each dataset is evenly divided into two subsets: one for member data used to train the teacher model (EDM) and the other for non-member data (detailed EDM training setup is in App. A). The teacher model generates 100,000 synthetic samples for the distillation of the student model, ensuring that the student model never accesses the original training data of the teacher model. We construct an auxiliary non-member dataset by randomly sampling 15,000 data points from the non-member data of FFHQ and CIFAR10, with 5,000 points used for deep-kernel training (Alg. 1) and 10,000 for candidate dataset detection (Alg. 2). For AFHQv2, we sample 3,000 non-member data points, allocating 1,500 for kernel training and 1,500 for candidate detection. To ensure fairness, we randomly discard 15,000 member data points (3,000 for AFHQv2).
To evaluate D-MIA under varying proportions of member data in the candidate datasets, we create candidate datasets with 100%, 50%, and 30% member data. During detection, we randomly sample 5,000 data points (1,500 for AFHQv2) based on the specified member ratios to construct positive candidate datasets. Additionally, we construct a negative candidate dataset consisting entirely of non-member data to assess whether it can be distinguished from the positive datasets. Similar to the baseline setting, we perform 50 rounds of sampling and detection to verify the attack accuracy of D-MIA.
Implementation details of D-MIA The network architecture of the deep-kernel MMD follows the design proposed by Liu et al. (2020). The training parameters (e.g., bandwidth, learning rate, and epochs) used for attacking different models with various training datasets are detailed in Tab. 5.
Table 5: Deep-kernel training configurations for distillation models (DI and DMD) across different datasets. “Bandwidth” denotes the kernel bandwidth used in the deep-kernel MMD loss; “Epoch” indicates the total number of training iterations of deep-kernel; “MMD learning rate” refers to the learning rate of the deep-kernel MMD training; “H” represents the number of hidden features or layers used in the feature extractor network; and “x_out” is the output dimensionality of the feature extractor network.
Model Dataset Bandwidth Epoch MMD learning rate H x_out
DI CIFAR10 0.1 400 0.000001 450 35
FFHQ 0.4 300 0.000001 450 50
AFHQv2 0.1 400 0.000001 450 35
DMD CIFAR10 0.0025 300 0.0000001 250 20
FFHQ 0.4 300 0.000001 450 50
AFHQv2 0.1 400 0.000001 450 35
B.4 Key algorithms in D-MIA
This section details the three key steps in D-MIA, each executing a specific algorithm: Deep-Kernel Training (Alg. 1), Detecting The Candidate Dataset (Alg. 2), and Ensembling Multiple Kernels. (Alg. 3). Fig. 4 overviews the framework pipeline.
Algorithm 1 Deep-Kernel Training
1:  Input: non-member dataset ${\mathcal{D}}^{\rm non}$; one-step generative model $G_{S}$ and encoder $G_{e}$ of $G_{S}$;          standard deviation $\sigma$ of additive Gaussian noise; learning rate $\eta$; epochs $E$
2:  $S_{g}\leftarrow\{G_{S}(z_{i})\mid z_{i}\sim\mathcal{N}(0,I),\ i=1,\dots,N\}$
3:  $S_{g,\mathrm{noisy}}\leftarrow\{s+\epsilon\mid s\in S_{g},\ \epsilon\sim\mathcal{N}(0,\sigma^{2}I)\}$
4:  $S_{a,\mathrm{noisy}}\leftarrow\{a+\epsilon\mid a\in{\mathcal{D}}^{\rm non},\ \epsilon\sim\mathcal{N}(0,\sigma^{2}I)\}$
5:  $S_{g-e}\leftarrow\{G_{e}(s)\mid s\in S_{g,\mathrm{noisy}}\}$
6:  $S_{a-e}\leftarrow\{G_{e}(a)\mid a\in S_{a,\mathrm{noisy}}\}$
7:  Sample mini-batch ${\mathcal{B}}^{\rm non}\subset S_{a-e}$
8:  Sample mini-batch ~$\tilde{{\mathcal{B}}}^{\rm mem}\subset S_{g-e}$
9:  Sample mini-batch ${\mathcal{B}}^{\rm anc}\subset S_{g-e}$ such that ${\mathcal{B}}^{\rm anc}\cap\tilde{{\mathcal{B}}}^{\rm mem}=\emptyset$
10:  for $\text{epoch}=1$ to $E$ do
11:     $M_{1}\leftarrow\widehat{\operatorname{MMD}}^{2}_{u}({\mathcal{B}}^{\rm non},{\mathcal{B}}^{\rm anc},k_{\omega})$
12:     $M_{2}\leftarrow\widehat{\operatorname{MMD}}^{2}_{u}(\tilde{{\mathcal{B}}}^{\rm mem},{\mathcal{B}}^{\rm anc},k_{\omega})$
13:     $l\leftarrow M_{1}-M_{2}$
14:     $\omega\leftarrow\text{Adam}(\omega,\nabla l,\eta)$
15:  end for
16:  Output: trained kernel $k_{\omega}$; anchor features ${\mathcal{B}}^{\rm anc}$
Algorithm 2 Detecting the Candidate Dataset
1:  Input: non-member dataset ${\mathcal{D}}^{\rm non}$; candidate dataset ${\mathcal{D}}^{\rm can}$; anchor features ${\mathcal{B}}^{\rm anc}$; encoder $G_{e}$ of $G_{S}$;          Gaussian noise std. $\sigma$; number of repetitions $T$; kernel function $k_{\omega}$
2:  $S_{c,\mathrm{noisy}}\leftarrow\{c+\epsilon\mid c\in{\mathcal{D}}^{\rm can},\ \epsilon\sim\mathcal{N}(0,\sigma^{2}I)\}$
3:  $S_{a,\mathrm{noisy}}\leftarrow\{a+\epsilon\mid a\in{\mathcal{D}}^{\rm non},\ \epsilon\sim\mathcal{N}(0,\sigma^{2}I)\}$
4:  $S_{c-e}\leftarrow\{G_{e}(c)\mid c\in S_{c,\mathrm{noisy}}\}$
5:  $S_{a-e}\leftarrow\{G_{e}(a)\mid a\in S_{a,\mathrm{noisy}}\}$
6:  Sample mini-batch ${\mathcal{B}}^{\rm can}\subset S_{c-e}$
7:  Sample mini-batch ${\mathcal{B}}^{\rm non}\subset S_{a-e}$
8:  for $t=1$ to $T$ do
9:     $M^{t}_{1}\leftarrow\widehat{\operatorname{MMD}}^{2}_{u}({\mathcal{B}}^{\rm can},{\mathcal{B}}^{\rm anc},k_{\omega})$
10:     $M^{t}_{2}\leftarrow\widehat{\operatorname{MMD}}^{2}_{u}({\mathcal{B}}^{\rm non},{\mathcal{B}}^{\rm anc},k_{\omega})$
11:     ${\mathbb{I}}^{(t)}\leftarrow\mathbbm{1}(M^{t}_{1}<M^{t}_{2})$
12:  end for
13:  $p^{\rm mem}\leftarrow\frac{1}{T}\sum_{t=1}^{T}{\mathbb{I}}^{(t)}$
14:  Output: $p^{\rm mem}$
Algorithm 3 Ensembling Multiple Kernels
1:  Input: anchor features ${\mathcal{B}}^{\rm anc}$; non-member dataset ${\mathcal{D}}^{\rm non}$; candidate dataset ${\mathcal{D}}^{\rm can}$;          one-step generative model $G_{S}$ and encoder $G_{e}$; number of iterations $h$; threshold $\tau$;          Gaussian noise std. $\sigma$; test repetitions $T$; learning rate $\eta$; epochs $E$
2:  Initialize kernel set $K\leftarrow\emptyset$, prediction set $R\leftarrow\emptyset$
3:  for $i=1$ to $h$ do
4:     Train kernel $k^{i}_{\omega}$ using Algorithm 1
5:     Compute $p^{\rm mem}_{(i)}$ using Algorithm 2
6:     $K\leftarrow K\cup\{k^{i}_{\omega}\}$
7:     $R\leftarrow R\cup\{p^{\rm mem}_{(i)}\}$
8:  end for
9:  $\bar{p}^{\rm mem}\leftarrow\frac{1}{h}\sum_{i=1}^{h}p^{\rm mem}_{(i)}$
10:  if $\bar{p}^{\rm mem}\geq\tau$ then
[…]
12:  else
[…]
14:  end if
[…]
B.5 D-MIA’s reliance on auxiliary non-member and candidate dataset sizes
In D-MIA attacks, the attacker requires a certain amount of non-member data for auxiliary training and testing. Additionally, the candidate dataset being evaluated must have a sufficient size to obtain accurate distributional information. Therefore, we evaluate the performance of D-MIA on CIFAR10 models for DMD and DI under different auxiliary non-member dataset sizes and candidate dataset sizes. We evaluated three settings for auxiliary and candidate dataset sizes: auxiliary dataset sizes of 15,000, 9,000, 6,000, and 3000 paired with candidate dataset sizes of 5,000, 3,000, 2,000, and 1000, respectively. Half of the auxiliary dataset was used to train the deep-kernel, while the other half was used to support attacks on the candidate dataset. Positive samples were drawn from member data corresponding to the candidate dataset size, and negative samples were drawn from non-member data of the same size. Following the previous evaluation, 50 positive and 50 negative samples were constructed, and D-MIA was applied to distinguish between them.
Appendix C Additional Experimental Results
We conducted a series of experiments to evaluate the effectiveness of different I-MIA methods on various generative models. Specifically, we extracted half of the data from the CIFAR10, FFHQ, and AFHQv2 datasets to train three EDM generative models, and then used the data generated by EDM to train DMD and Diff-Instruc. Finally, we applied four state-of-the-art MIA techniques—GAN-Leak, SecMI, ReDiffuse, and GSA—to attack these models. The ASR and AUC results are presented in Tab. 6. The TPR values at FPR = 0.05 results are presented in Tab. 7.
Table 6: the ASR and AUC results of various membership inference attack methods across different generative models and datasets. The table compares four attack methods—GAN-leak, SecMI, ReDiffuse, and GSA—on three generative models: EDM, DMD, and Diff-Instruc, evaluated on CIFAR-10, FFHQ, and AFHQv2 datasets.
Model/Dataset GAN-leak SecMI Rediffuse GSA
ASR AUC ASR AUC ASR AUC ASR AUC
EDM/CIFAR10 0.536 $\pm$ .005 0.523 $\pm$ .011 0.588 $\pm$ .004 0.601 $\pm$ .021 0.579 $\pm$ .002 0.603 $\pm$ .004 0.622 $\pm$ .008 0.626 $\pm$ .004
EDM/ffhq 0.524 $\pm$ .008 0.518 $\pm$ .018 0.551 $\pm$ .009 0.564 $\pm$ .011 0.541 $\pm$ .005 0.553 $\pm$ .005 0.662 $\pm$ .006 0.654 $\pm$ .003
EDM/afhqv 0.543 $\pm$ .004 0.532 $\pm$ .009 0.604 $\pm$ .005 0.622 $\pm$ .013 0.604 $\pm$ .005 0.644 $\pm$ .006 0.906 $\pm$ .004 0.908 $\pm$ .001
DMD/CIFAR10 0.497 $\pm$ .012 0.508 $\pm$ .011 0.520 $\pm$ .018 0.516 $\pm$ .020 0.514 $\pm$ .008 0.509 $\pm$ .013 0.512 $\pm$ .003 0.502 $\pm$ .001
DMD/ffhq 0.502 $\pm$ .019 0.498 $\pm$ .021 0.515 $\pm$ .021 0.502 $\pm$ .037 0.507 $\pm$ .004 0.504 $\pm$ .008 0.525 $\pm$ .002 0.505 $\pm$ .001
DMD/afhqv 0.512 $\pm$ .009 0.515 $\pm$ .032 0.525 $\pm$ .007 0.513 $\pm$ .007 0.521 $\pm$ .007 0.524 $\pm$ .004 0.532 $\pm$ .004 0.523 $\pm$ .003
Diff-Instruc/CIFAR10 0.502 $\pm$ .005 0.497 $\pm$ .003 0.507 $\pm$ .004 0.501 $\pm$ .009 0.514 $\pm$ .004 0.511 $\pm$ .007 0.503 $\pm$ .001 0.503 $\pm$ .001
Diff-Instruc/ffhq 0.493 $\pm$ .002 0.503 $\pm$ .005 0.514 $\pm$ .008 0.509 $\pm$ .008 0.509 $\pm$ .002 0.509 $\pm$ .004 0.501 $\pm$ .002 0.511 $\pm$ .002
Diff-Instruc/afhqv 0.501 $\pm$ .009 0.502 $\pm$ .006 0.504 $\pm$ .005 0.504 $\pm$ .008 0.513 $\pm$ .003 0.506 $\pm$ .005 0.511 $\pm$ .005 0.515 $\pm$ .002
Figure 5: Distribution analysis of D-MIA outputs across different member/non-member ratios within the candidate sets. Results are shown for distilled models against CIFAR10 (a, c) and FFHQ (b, d), where subfigures (a, b) report the results of DMD, while subfigures (c, d) present the results of Diff-Instruct.
Table 7: TPR values at FPR = 0.05 for three MIA methods (SecMI, ReDiffuse, GSA) across different generative models and datasets.
Model Dataset SecMI ReDiffuse GSA
EDM CIFAR10 0.07 0.06 0.08
FFHQ 0.09 0.07 0.09
AFHQv2 0.11 0.13 0.45
DMD CIFAR10 0.05 0.05 0.05
FFHQ 0.04 0.05 0.05
AFHQv2 0.06 0.07 0.06
Diff-Instruct CIFAR10 0.04 0.04 0.05
FFHQ 0.05 0.05 0.05
AFHQv2 0.06 0.05 0.05
Table 8: True positive rates (TPR) of D-MIA, SecMI, and Rediffuse at a fixed false positive rate (FPR) of 0.05 under varying member proportions (30%, 50%, 100%) are reported across three datasets—CIFAR10, FFHQ, and AFHQv2—for both DMD and Diff-Instruct (DI) models.
Model Dataset Member Ratio DGG-MIA SecMI Rediffuse
DMD CIFAR10 100% 0.88 0.12 0.10
50% 1.00 0.00 0.00
30% 0.96 0.00 0.02
FFHQ 100% 0.96 0.04 0.04
50% 0.98 0.02 0.04
30% 1.00 0.02 0.08
AFHQV 100% 1.00 0.16 0.26
50% 1.00 0.08 0.12
30% 1.00 0.10 0.16
DI CIFAR10 100% 1.00 0.02 0.04
50% 1.00 0.02 0.10
30% 1.00 0.02 0.02
FFHQ 100% 1.00 0.12 0.28
50% 1.00 0.04 0.10
30% 1.00 0.08 0.14
AFHQV 100% 1.00 0.04 0.08
50% 1.00 0.00 0.08
30% 1.00 0.04 0.12